An open pipe 30cm long and a closed pipe 23cm long both of the same diameter are each sending their first overtone are in unison. the end correction of these pipes is

An  open pipe 30cm long and a closed pipe 23cm long both of the same diameter are each sending their first overtone are in unison. the end correction of these pipes is

Grade:12

2 Answers

Ashish Garg
5 years ago
Suppose v Is the velocity of light and n is Me frequency of the note.When tho open lalos io wunging its first overtone. the length of the pipe plus end correction = 1 = =30 1.2cSince there are Wo end cone-When Me closed pipe is sound- ing its first overtone 4' = c=23+c or,n = 23+c •00 From 01 en (0) 23 +0 = 1=00 1.28)92+4c = 90+ 8c c = 1 cm
Radha Shukla
5 years ago
For cop v/4L+e and for oop v/2(L+e) equate them do remember to put length correctly and e=0.3d since they have same diameter so e will be same for both